8. Calculating inflation using a simple price index Consider a fictional price index, the College Student Price Index (CSPI), based on a typical college students annual purchases. Suppose the following table shows information on the market basket for the CSPI and the prices of each of the goods in 2010, 2011, and 2012 The cost of each item in the basket and the total cost of the basket are shown for 2010 Perform these same calculations for 2011 and 2012, and enter the results in the following table 2010 2011 2012 Price Cost Price Cost Price Cost Notebooks Calculators Large coffees Energy drinks Textbooks Total cost Price index Quantity in Basket (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) 40 104 600 375 960 2,079 10 30 75 600 150 720 1,575 100 3 30 80 600 300 880 4 75 80 104 300 75 8 4 5 90 110 120 1,890 120Suppose the base year for this price index is 2010. In the last row of the table, calculate and enter the value of the CSPI for the remaining years. Between 2010 and 2011, the CSPI increased by 20% . Between 2011 and 2012, the CSPI increased by which of the following, if true, would illustrate why price indexes such as the CSPI might overstate inflation in the cost of going to college? Check all that apply. A new type of personal transporter, which made it easier to get around places like university campuses, became available for purchase. Professors required each student to buy 10 notebooks, regardless of the price. The quality of textbooks increased dramatically from 2010 to 2012, with textbook companies bundling new online study aids with their books As the price of energy drinks increased relative to the price of coffee between 2010 and 2012, students decreased their consumption of energy drinks and increased their consumption of coffee.Please check and correct my work and answer the rest. Thank you :)
Answer #1

Cost in 2011 = price in 2011*respective quantities
Notebooks = 3*10 = 30
Calculators = 80*1 = 80
Large coffees = 2*300 = 600
Energy drinks = 4*75 = 300
Textbooks = 110*8 = 880
Total cost of basket in 2011 = 30 + 80 + 600 + 300 + 880 = 1,890
Price index in 2011 = (Total cost of basket in 2011/Total cost of basket in 2010)*100 = (1,890/1,575)*100 = 120

Cost in 2012 = price in 2012*respective quantities
Notebooks = 4*10 = 40
Calculators = 104*1 = 104
Large coffees = 2*300 = 600
Energy drinks = 5*75 = 375
Textbooks = 120*8 = 960
Total cost of basket in 2011 = 40 + 104 + 600 + 375 + 960 = 2,079
Price index in 2012 = (Total cost of basket in 2012/Total cost of basket in 2010)*100 = (2079/1,575)*100 = 132

Between 2010 and 2011, increase in CSPI = [(Price index in 2011-Price index in 2010)/Price index in 2010]*100
= [(120-100)/100]*100 = 20%
Between 2011 and 2012, increase in CSPI = [(Price index in 2012-Price index in 2011)/Price index in 2011]*100
= [(132-120)/120]*100 = 10%

Option A, C, and D will overstate inflation.